OpenStudy (anonymous):

v^2+3v-40=0 is the answer 0?

OpenStudy (anonymous):

v=0?

OpenStudy (anonymous):

no, x=-8 or 5 \[v^{2}+3v-40=0\] \[v^{2}-5v+8v-40=0\] \[v\left( v-5 \right)+8\left( v-5 \right)=0\] \[\left( v+8 \right)\left( v-5 \right)=0\] v+8=0, v-5=0 v=-8 or v=5

OpenStudy (anonymous):

so would the anser be v=-8,5?

OpenStudy (anonymous):

yes